40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 91 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  12516   Fri Sep 23 01:09:04 2016 ericqUpdateSUSETMX hysteresis test

I had hoped to do some ALS work, but I realized too late that we loaned our HP analyzer to Andrew. I decided instead to do some ETMX testing. 

I have a script running that'll misalign both ETMs and back by about 0.5mrad with half hour rests in between. It'll be done around 6AM.

  12515   Thu Sep 22 22:52:08 2016 ericqUpdateGeneralDamping found to be on

Just a heads up, it looks like the damping came on at around 8:30pm. Not sure why. 

  12514   Thu Sep 22 20:18:27 2016 LydiaUpdateGeneralAcromag Progress

We moved the Acromag and its power supply to the X end, where we connected it to the diagnostic output of the NPRO controller. We renamed the channels to be descriptive of the pin outputs as described in the laser manual. We were able to recover readouts similar to those we found with a multimeter. 

We should figure out how to set up the channels on the front end machines: right now they are accessed through a tmux session running on pianosa. Once we are confident in the operation, we will make a box to contain the Acromag and wire connections and move the setup to connect to the PSL controller. 

  12513   Thu Sep 22 20:01:47 2016 LydiaUpdateGeneralITMX freed, all optics kicked

Rana came by and freed ITMX again. I think it shouldn't be a problem for me to free it if it happens again. 

In hopes of getting better SNR on the free swing spectra, we kicked all optics at around 7pm. The damping should come back on a little after midnight. ITMX did not get stuck after this kick. 

  12512   Thu Sep 22 15:44:21 2016 SteveUpdateGeneralITMX magnets are stucked again

 

 

Attachment 1: StuckAgain.png
StuckAgain.png
  12511   Wed Sep 21 09:04:57 2016 SteveUpdateGeneral8 hours recovery progress

Good 8 hours

Quote:

The misalignment wasn't as bad as I had intially feared; the spot was indeed pretty high on ETMX at first. Both transmon QPDs did need a reasonable amount of steering to center once the dither had centered the beam spots on the optics.

Arms, PRMI and DRMI have all been locked and dither aligned. All oplevs and transmon QPDs have been centered. All AS and REFL photodiodes have been centered. 

Green TM00 modes are seen in each arm; I'll do ALS recovery tomorrow. 

 

Attachment 1: 8hrs.png
8hrs.png
  12510   Wed Sep 21 01:08:02 2016 ericqUpdateGeneralFurther recovery progress

The misalignment wasn't as bad as I had intially feared; the spot was indeed pretty high on ETMX at first. Both transmon QPDs did need a reasonable amount of steering to center once the dither had centered the beam spots on the optics.

Arms, PRMI and DRMI have all been locked and dither aligned. All oplevs and transmon QPDs have been centered. All AS and REFL photodiodes have been centered. 

Green TM00 modes are seen in each arm; I'll do ALS recovery tomorrow. 

  12509   Tue Sep 20 17:04:46 2016 SteveUpdateElectronicsREF33

REF33 was removed for taking picture of the bare C30362 InGaAs photodiode per Rana's request. All other rf photodiodes have their glass cover on.

Note: it is back to it's place but this pd will need alignment!

The small steering mirror was completly lose before it was removed.

Attachment 1: A005_-_20160920_135529_-_Shortcut.lnk.bmp
  12508   Tue Sep 20 10:45:06 2016 ranaUpdateGeneralFurther recovery progress

Rana suspicious. We had arms locked before pumpdown with beams on Transmon PDs. If they're off now, must be beams are far off on the mirrors. Try A2L to estimate spot positions before walkin the beams too far.

  12507   Mon Sep 19 22:03:10 2016 ericqUpdateGeneralFurther recovery progress

[ericq, Lydia, Teng]

Brief summary of this afternoon's activities:

  • PMC alignment adjusted (Transmission of 0.74)
  • IMC locked, hand aligned. Tranmission slightly over 15k. Measured spot positions to be all under 2mm.
  • Set DC offsets of MC2 Trans + WFS1 + WFS2 (WFS2 DC offsets had wandered so much that DC "centered" left some quadrants almost totally dark)
  • Set demod offsets of WFS1+WFS2
    • Note to self: WFS script area is a mess. I can never remember which scripts are the right ones to run. I should clean this up
  • WFS loops activated, tested. All clear.
  • Locked Yarm, dither aligned. Transmission 0.8
  • Moved BS to center ITMY reflection on AS camera
  • Misaligned ETMY, aligned PRM to make a flashing PRY AS beam. REFL camera spot confirmed to be on the screen, which is nice
  • Wandered ITMX around until its AS spot was found. ITMX OSEMs not too far from their half max. (todo: update with numbers)
  • Wandered SRM around until full DRMI flashes seen
  • Centered all vertex oplevs
  • Made a brief attempt at locking X arm, could only get some crazy high order mode to lock. BS and ITMX alignments have changed substantially from the in-air locks, so probably need to adjust ETMX much more.

Addendum: I had a suspicion that the alignment had moved so much, we were missing the TRX PDs. I misaligned the Y arm, and used AS110 as a proxy for X arm power, as we've done in the past for this kind of thing. Indeed, I could maximize the signal and lock a TM00 mode. Both the high gain PD and QPD in the TRX path are totally dark. This needs realignment on the end table.

  12506   Mon Sep 19 13:57:21 2016 ranaUpdateGeneralnever post EPS files in the ELOG. Ever.

http://tex.stackexchange.com/questions/2092/which-figure-type-to-use-pdf-or-eps

  12505   Mon Sep 19 13:25:03 2016 TengUpdateElectronicsSatellite Amplifier

 

In order to figure out the difference betweent simulated result and measurement, I tried to measuren the electronic noise by following ways as show in attachment 1

1.measure from the satellite box by SR785 at ETMY ,calibrate to counts by divide by 3267.8. while at that conditin, the set up is in suspension.

2. measure after ADC by diagnostics test tools, with set up on table in history and on uspension currently.

3. use the caculated butterfly channel.

the results are shown in attachmemt 2. The overall nosie level are still much higher than simulation.

 

 

Quote:

 

If we have some data with one of the optics clamped and the open light hitting the PD, or with the OSEMs removed and sitting on the table, that would be useful for evaluating the end-to-end noise of the OSEM circuit. It seems like we probably have that due to the vent work, so please post the times here if you have them.

The ETMX OSEMs have been attached to its Satellite box and plugged in for the last 10 days or so, with the PD exposed to the unobstructed LED. I pulled the spectrum of one of the sensors (mean detrended, I assume this takes care of removing the DC value?). The DQed channels claim to record um (the raw ADC counts are multiplied by a conversion factor of 0.36). For comparison, re-converted the y-axis for the measured curve to counts, and multiplied the total noise curve from the LISO simulation by a factor of 3267.8cts/V (2^16cts/20V) so the Y axis is noise in units of counts/rtHz. At 1Hz, there is more than an order of magnitude difference between the simulation and the measurement which makes me suspect my y-axis conversion, but I think I've done this correctly. Can such a large discrepancy be solely due to thick film resistors?

 

  12504   Mon Sep 19 11:11:43 2016 ericqUpdateSUSAlignment status

[ericq, Steve]

ITMX is free, OSEM signals all rougly centered. 


This was accomplished by rocking the static alignment (i.e. slow controls) pitch and yaw offsets until the optic broke free. This took a few volts back and forth. At this point, I tried to find a point where the optic seemed to freely swing, and hopefully have signals in all 5 OSEMS. It seemed to be free sometimes but mostly settling into two different stationary states. I realized that it was becoming torqued enough in pitch to be leaning on the top-front or top-back EQ stops. So, I slowly adjusted the pitch from one of these states until it seemed to be swinging a bit on the camera, and three OSEM signals were showing real motion. Then, I slowly adjusted the pitch and yaw alignments to get all OSEMS signals roughly centered at half of their max voltage.

  12503   Sun Sep 18 16:18:05 2016 ranaUpdateSUSAlignment status

susaux is responsible for turning on/off the inputs to the coil driver, but not the actual damping loops. So rebooting susaux only does the same as turning the watchdogs on/off so it shouldn't be a big issue.

Both before and after would be good. We want to see how much bias and how much voltage from the front ends were applied. l1susaux could have put in a huge bias, but NOT a huge force from the damping loops. But I've never seen it put in a huge bias and there's no way to prevent this anyway without disconnecting cables.

I think its much more likely that its a little stuck due to static charge on the rubber EQ stop tips and that we can shake it lose with the damping loops.

  12502   Sat Sep 17 16:51:01 2016 LydiaUpdateSUSAlignment status

Here's the timeseries plots. I've zoomed in to right after the problem- did you want before? We pretty much know what happened: c1susaux was restarted from the crate but the damping was on, so as soon as the machine came back online the damping loops sent a huge signal to the coils. (Also, it seems to be down again. Now we know what to do first before keying the crate.) It seems like both right side magnets are stuck, and this could probably be fixed by moving the yaw slider. Steve advised that we wait for an experienced hand to do so. 

Quote:

All is not lost. I've stuck and unstuck optics around a half dozen times. Can you please post the zoomed in time series (not trend) from around the time it got stuck? Sometimes the bias sliders have to be toggles to make the bias correct. From the OSEM trend it seems like it got a large Yaw bias. May also try to reseat the satellite box cables and the cable from the coil driver to the cable breakout board in the back of the rack.

 

Attachment 1: Screenshot_from_2016-09-17_16-45-00.png
Screenshot_from_2016-09-17_16-45-00.png
  12501   Sat Sep 17 02:00:23 2016 ranaUpdateSUSAlignment status

All is not lost. I've stuck and unstuck optics around a half dozen times. Can you please post the zoomed in time series (not trend) from around the time it got stuck? Sometimes the bias sliders have to be toggles to make the bias correct. From the OSEM trend it seems like it got a large Yaw bias. May also try to reseat the satellite box cables and the cable from the coil driver to the cable breakout board in the back of the rack.

  12500   Fri Sep 16 19:48:52 2016 LydiaUpdateGeneralAlignment status

Today the Y arm was locking fine. The alignment had drifted somewhat so I ran the dither and TRY returned to ~0.8. However, the mode cleaner has been somewhat unstable. It locked many times but usually for only a few minutes. Maybe the alignment or autolocker needs to be adjusted, but I didn't change anything other than playing with the gain sliders (which didn't seem to make it either better or worse).

ITMX is still stuck.

  12499   Fri Sep 16 19:14:27 2016 LydiaUpdateSUSDiagonalization

[Lydia, Teng]

We built matrices for ITMY and ETMY by driving one degree of freedom at a time with awggui, while the damping was on. These have been applied to the damping loops.

  • Each segment of data is 1000s long and each dof was driven at 0.25 Hz.
  • These matrices are much closer to the ideal matrix and have no wrong signs. We believe they represent the relative sensitivity of the OSEMs to the degrees of freedom much more accurately. This is because the free swinging modes are not actually pitch, yaw, etc, but some linear combination of these. However, the damping actuates on pitch, yaw, etc. So we should isolate the degrees of freedom by driving them one at a time instead of just looking at free swinging peaks.
    • Attachment 1: An example of the dof spectra, calculated using the default input matrix, when ETMY YAW was driven at 0.25 Hz.
    • Attachment 2: The same OSEM sensor data, with the dofs calculated using the matrix found from this data. There is still a significant peak in pitch, but the other dofs are significantly suppressed.
    • Attahcment 3: The same data again, but the dofs are measured with the input matrix calculated by the free swinging data. This achieves much less suppression than the new matrix. Obviously this is not exactly a fair comparison because the new matrix was generated with this data, but the method of measuring OSEM responses by driving peaks has a much close relationship between what it measured (the OSEM response), and how the matrix is used (by damping loops which drive the coils in much the same way as awggui).
  • The phase problems seem to be mostly solved. Both Y arm test masses have some phase warnings, but they mostly occur with side. This can happen because the ideal matrix elements are 0, so the real parts are small. If there is no strong coupling then there is no reason to expect the background spectrum to be in phase with the peak. Other phase differences are small; most less than 5 degrees, a couple between 5 and 10 degrees. This may still merit further investiagtion.
  • Comparing the damping results for ITMY with the old (based on free swinging data) and new (based on driven data), we see the 1Hz peak suppressed by ~35% and the noise above 1Hz generally suppressed by ~25-30% . There is, however, significantly more movement between 0.5 and 1 Hz, maybe because the fundamental physical modes are not being directly measured and suppressed. Overall this seems like an improvement.

GPS times:

ITMY

Pitch:1158085097 Yaw: 1158086537 Pos: 1158089237 Side: 1158087977

ETMY

Pitch: 1158095897 Yaw: 1158097577 Pos: 1158099377 Side: 1158100817

Attachment 1: ETMY_yawdrivedefault.png
ETMY_yawdrivedefault.png
Attachment 2: ETMY_yawdrivenew.png
ETMY_yawdrivenew.png
Attachment 3: ETMY_yawdriveold.png
ETMY_yawdriveold.png
Attachment 4: 57.png
57.png
  12498   Fri Sep 16 14:15:28 2016 steveUpdatePEMpartical counts

South end flow bench and both clean room assembly flow benches measured zero counts for 0.3 and 0.5 micron size particales.

The counting efficiency of 0.5 micron is 100%

 

  location

 

 

        0.3 micron particles / cf min

 

 

      0.5 micron particles / cf min

 

 

 counter

 MET

 ONE

 

       effect
 ITMY table  35,000  3,000   #3  
 ETMY table  35,000  3,000   #3  
 ITMX table  24,000  2,400   #3  PSL HEPA at 50V
 ITMX wall mounted  -  2,500   #1  
 ETMX work bench    5,000     600   #3   flow bench on  

 

The PSL HEPA performance was measured at the center of the table with MET ONE #3

 Voltage Variac particles /cf min   particles /cf min
 particle size  0.3 micron  0.5 micron
 60  0  0
 50  10  0
 40  1,400  110
 20  1,400  100
Attachment 1: 80days.png
80days.png
  12497   Thu Sep 15 18:37:20 2016 LydiaUpdateSUSDiagonalization

[Teng, Lydia]

  • We fixed the 60Hz filter on ITMY. This improved the phase problems somewhat but one coil (UL) is still about 12 degrees out of phase compared to the others for all the dofs. Is there some other place where a filter coule be applied to just one coil sensor? I pressed the "Load coefficients" button for UL, so maybe that will have helped.
  • We want to interpret the coil signals to have an accurate measurement of each dof. This means what the input matrix should describe is the dependence of each dof on the OSEM signals, which is found by inverting the matrix which describes the sensitivity of each OSEM to changes in that degree of freedom.
    • We looked at the spectra of the individual coils for ITMY and ETMY (See attachment 1 & 2). The coupling between some coils and applicable resonance peaks is very weak (~0.1 times the sensitivity of the other coils).
    • However, when a certain degree of freedom, e.g. pitch, is deliberately driven using awggui, the response of the ITMY coils is clear on the StripTool and is about the same magnitude for all of the face OSEMS. So, it seems like the diagonalization script does not always succeed at measuring the relative sensitivity of the OSEMs to the degrees of freedom.
    • This may be because the fundamental swing modes experienced by the free swinging pendulum are not the same as what we measure as pitch, yaw, etc. This could be possible if the wire tension is not the same on both sides. For ITMY, the spectra imply that the funamdental frequencies are actually at some linear combinations of pitch and yaw, swinging about a diagonal axis that results in a much weaker response for some of the OSEMS. Calling these peaks pitch and yaw may be inaccurate. Certainly they do not indicate the true relative sensitivity of the coils.
    • We propose an alternate approach to measuring this sensitivity: drive one dof at a time with awggui, take a spectrum (less resolution is ok because we already know the drive frequency), and measure the sensing matrix values for that dof the same way as before, but using a spectral peak that decribes motion that we know is purely pitch. Repeat this for all 4 dofs that we can actuate on, then compile these results into a sensing matrix and take the inverse.
Attachment 1: ETMY_osemspec.png
ETMY_osemspec.png
Attachment 2: ITMY_osemspec.png
ITMY_osemspec.png
  12496   Thu Sep 15 10:36:32 2016 SteveUpdateVACpumpdown 79 day 3

IFO pressure 3.7e-5 Torr at new cold cathode InstruTech - Hornet

New items in vacuum:

1, ETMX sus tower with new baked sus wire, EP30-2 epoxied magnets, same at different locations also........ ..........  and 2 ruby wirestandoffs.

2, First Contact cleaned arm test masses only. This technic was a 1st time use in our vacuum system.

3, 50 mm ID green glass baffles at the ends

4, witness mirrors at ETMX and ITMY (old oplev mirrors) We observed a very dusty system: sides of optics, towers and tables were wiped. Hepa tents used at Y arm and BS

5, new pirani, cc gauge and 1.5" right angle valve

 

Quote:

The pumpdown started at 9-12-2016

The IFO is at 5e-5 Torr vacuum normal after 73 day at atm.

ITMX needed to be freed and ETMY-UL is still misbehaving occasionally.

New pirani   and cold cathode gauges added at this vent. They were baked at 100 C for 6 hrs under vacuum.

Go to the Vac Rack to read IFO pressure from the gauge itself when Vac. Monitor is blank as it is now !

 

Quote:

P1 IFO pressure is 1 mTorr, valve configuration: vacuum normal, annulosses are pumped, RGA is not pumped.

THANKS to Chris !

The shutters can be opened with high power.

No communication error message still exist.

I will reboot c1vac1 and c1vac2 to get gauge communication with medm tomorrow.

 

 

 

 

Attachment 1: vent78for73d.png
vent78for73d.png
  12495   Wed Sep 14 20:27:03 2016 LydiaUpdateSUSDiagonalization

Today the main optics were free swinging for several hours, so I attempted diagonalization in vacuum.

  • ITMY still has bad phases. I looked at the spectra for this and other optics, and it looks like the other optics have the 60Hz line notched out for all coils while ITMY only has it notched on the side coil. (Using C1:SUS-ITMY_SENSOR channels). Where is this controlled from, and could it be the source of the issue? 
    • I tried using a different coil as the "standard," with the other coils compared against it in tfestimate. Default is UL, I tried UR and LL. The phase problems were still present for ITMY, but the script was still working fine for other optics.
    • The phase difference between coils is different for different start times.
    • A short segment of the time series for ITMY shows significantly more high frequency noise than for other optics at the same time.
  • The ETMY matrix for vacuum has the wrong sign for UL coupling to pitch! The diagonalization results look OK on the graph, but the butterfly mode still has small peaks (See attachment 1). When the individual coil spectra are plotted, the angular degrees of freedom show very weak coupling for UL to pitch, and LL to yaw. We initially replaced the matrix on the MEDM screen with the one generated by the script. After realizing this, the PIT row was changed to 1 1 -1 -1 0, but the effectiveness of the damping on the locked transmission fluctuations was about the same both ways.
Attachment 1: ETMY_diag.png
ETMY_diag.png
  12494   Wed Sep 14 20:05:32 2016 LydiaUpdateGeneralITMX magnets are stucked

When I restarted c1susaux yesterday, I didn't know that I needed to disable the coil outputs first. So when it came back online, it attempted to damp all the vertex area optics and ITMX got stuck frown

We should make a note in the Computer Restart Procedures wiki page indicating the importance of disabling the coils before rebooting c1susaux, c1auxex, and c1auxey. Today c1auxey was rebooted properly without incident. If the slider values etc go back to their previous values on their own, is it necessary to do a BURT restore? I tried doing one for c1susaux today and there were some errors for ASC channels, but the alignment sliders went right back to the proper place after reboot yesterday.

Quote:

I believe that the UR and LR magnets are stuck. There was no earth quake at 16:18 yesterday. Something had to kick it into this position. See 4days plot

Please advise  freeing details

 

  12493   Wed Sep 14 19:41:23 2016 JohannesUpdateGeneralPSL back to high power

Today's summary:

  • Replaced mirror in MC REFL path with R=10% BS and aligned beam on PD while still at low power
  • Replaced HR mirror in Transmon path at EY table with 50/50 BS. Alignment onto QPD not yet confirmed because we need IR from the YARM for it.
  • Put ND filters back on Transmon QPDs at both X and Y ends. For now I put all the filters on, for a combined OD of 1.6 at both ends (1.0 + 0.6 at YEND and 1.0 + 0.4 + 0.2 at XEND).
  • Put ND filter back on Transmon CCD on EY table.
  • Reverted MC autolocker to nominal, high power version.
  • Raised PSL output power back to nominal level by turning the waveplate. At the PSL shutter I measured a power of 1.03W. It occured to me too late that I realigned the PMC only afterwards and increased its transmission by a few percent, so I'll have to re-measure the actual PSL power.
  • MC is locked with its transmission back up to ~15,400 counts. The autolocker is not very good at obtaining the lock, as it seems to try to turn the VCO gain up too far and loses lock. The script probably needs a revision.
  • The YARM was pretty badly aligned. We used the green light to visually center the beam on the test masses AND had to go exploring with the TTs to see IR flashes in the first place. We got the YARM to lock to IR and were able to run the dither alignment. The maximum transmission we saw was on the order of 0.85. However, something strange is happening with the LSC control of the armlength. When the lock is engaged it drives PIT and YAW, which manifests itself in the OpLev signal and variable transmitted power on the TRY PDs. Osamu helped us diagnose this and was able to reduce the effect by tuning the POS gains to the individual ETMY OSEMS. The problem persisted even after using the new matrix diagonalization coefficients, we'll have to investigate this further and also take a look at the filters in the feedback path.
  • ITMX is still stuck and way out of alignment, so we couldn't even start with the green light in the XARM.
  12492   Wed Sep 14 09:44:09 2016 SteveUpdateVACpumpdown 79 completed

The pumpdown started at 9-12-2016

The IFO is at 5e-5 Torr vacuum normal after 73 day at atm.

ITMX needed to be freed and ETMY-UL is still misbehaving occasionally.

New pirani   and cold cathode gauges added at this vent. They were baked at 100 C for 6 hrs under vacuum.

Go to the Vac Rack to read IFO pressure from the gauge itself when Vac. Monitor is blank as it is now !

 

Quote:

P1 IFO pressure is 1 mTorr, valve configuration: vacuum normal, annulosses are pumped, RGA is not pumped.

THANKS to Chris !

The shutters can be opened with high power.

No communication error message still exist.

I will reboot c1vac1 and c1vac2 to get gauge communication with medm tomorrow.

 

 

 

Attachment 1: pd79.png
pd79.png
Attachment 2: pd79d2.jpg
pd79d2.jpg
Attachment 3: CCcg.jpg
CCcg.jpg
  12491   Wed Sep 14 09:33:54 2016 SteveUpdateGeneralITMX magnets are stucked

I believe that the UR and LR magnets are stuck. There was no earth quake at 16:18 yesterday. Something had to kick it into this position. See 4days plot

Please advise  freeing details.

Quote:

[Lydia,Teng]

Something strange happened to the ITMX osem reading around 4.pm. PDT as shown below.

Also the there was no response of the reading as we adjusted the PITCH and YAW. :(

Note that we restarted the slow machine: c1susaux,c1ausex this afternoon because of the unresponced interface.

 

 

 

Attachment 1: ITMXmagStuck.png
ITMXmagStuck.png
  12490   Tue Sep 13 19:18:43 2016 LydiaUpdateSUSDiagonalization in air

[Lydia, Teng]

We continued to work on the diagonalization scripts today and devised a way of choosing starting parameters that seems to work much better, and is easier to use, than tuning up to 15 parameters by hand per optic.

  • As before, the spectrum for each dof is estimated by using the "ideal" input matrix.
  • The starting guess for the peak frequency for each dof is the bin which achieves the maximum value of the spectrum between 0.4 and 1.5 Hz.
  • If another dof has a higher value at that frequency, the next highest peak is used. (Sometimes, for example, the peak in PIT at the POS frequency is stronger than the real POS peak!)
  • The peak height is initially guessed to be the spectrum value at the initial frequncy guess.
  • The width paramter Q can still be read from a file, but for all the times we tried, the peaks were found successfully if Q was initially guessed to be 300, so there might be no need to do this.
  • Spectra should still be examined to make sure the results make sense, and once we look at free swinging data in vacuum, we should compare the frequency results to the wiki values.
  • Reasonably good matrix values are saved to peakFit/inMats/1157630417. We got good diagonalization results for all but ITMY (see below). The values used for damping have not been overwritten.

We still noticed phase problems with ITMY, which appear to be preventing good diagonalization (See Attachment 1). Almost every degree of freedom has a significant imaginary part in the sensing matrix. We looked at the phases of the cross spectra in DDT and saw that indeed, the OSEM signals do not have the appropriate relative phases at the peak frequencies, especially in PIT and YAW (see Attachment 2: the phase at the peak is about 30 degrees when it should be 180). These phases are different for data takes ~24 hours apart, but are still wrong. We also looked at this information for ETMY and saw the correct behavior. We temporarily moved the pitch and yaw sliders for ITMY and looked at the OSEM response on a striptool, and the signals moved in the expected way. Can anyone suggest a reason why this would be happening? Is there another stretch of data (besides this past weekend) which would be good to compare to?

 

Attachment 1: ITMY_diag.jpg
ITMY_diag.jpg
Attachment 2: 38.png
38.png
  12489   Tue Sep 13 19:02:56 2016 TengUpdateGeneralITMX sensor

[Lydia,Teng]

Something strange happened to the ITMX osem reading around 4.pm. PDT as shown below.

Also the there was no response of the reading as we adjusted the PITCH and YAW. :(

Note that we restarted the slow machine: c1susaux,c1ausex this afternoon because of the unresponced interface.

 

 

Attachment 1: 47.png
47.png
Attachment 2: 34.png
34.png
  12488   Tue Sep 13 17:47:04 2016 SteveUpdateVACpumpdown 79 completed

P1 IFO pressure is 1 mTorr, valve configuration: vacuum normal, annulosses are pumped, RGA is off, not pumped.

THANKS to Chris !

The shutters can be opened with high power.

No communication error message still exist.

I will reboot c1vac1 and c1vac2 to get gauge communication with medm tomorrow.

 

Quote:

(Steve, Chris)

The pumpdown had stalled because of some ancient vacuum interlock code that prevented opening the valve V1 between the turbo pump and the main volume.

This interlock [0] compares the channels C1:Vac-P1_pressure and C1:Vac-PTP1_pressure, neither of which is functioning at the moment. The P1 channel apparently stopped reading sometime during the vent, and contained a value of ~700 torr, while the PTP1 channel contained 0. So the interlock code saw this huge apparent pressure difference and refused to move the valve.

To bypass this check, we used caput to enter a pressure of 0 for P1.

[0] /cvs/cds/caltech/state/from_luna/VacInterlock.st

 

  12487   Tue Sep 13 16:16:28 2016 ChrisOmnistructureVACvacuum interlock bypassed

(Steve, Chris)

The pumpdown had stalled because of some ancient vacuum interlock code that prevented opening the valve V1 between the turbo pump and the main volume.

This interlock [0] compares the channels C1:Vac-P1_pressure and C1:Vac-PTP1_pressure, neither of which is functioning at the moment. The P1 channel apparently stopped reading sometime during the vent, and contained a value of ~700 torr, while the PTP1 channel contained 0. So the interlock code saw this huge apparent pressure difference and refused to move the valve.

To bypass this check, we used caput to enter a pressure of 0 for P1.

[0] /cvs/cds/caltech/state/from_luna/VacInterlock.st

  12486   Tue Sep 13 11:00:59 2016 LydiaUpdateSUSETMY UL glitch returned

 

Quote:

For the ITMY, I squished together the cables which are in the 'Cable Interface Board' which lives in the rack. This thing takes the 64 pin IDC from the satellite module and converts it into 2 D-sub connectors to go to the PD whitening board and the coil driver board. Lets see if the ITMY OSEM glitches change character overnight.

Last night from 8:30 pm to 8:30 am PDT, ETMY UL signal was glitchy again. As of now it seems to have quieted back down, but we pushed on the cables on the board at the Y end to hopefully prevent it from coming back. After doing so it still seems to be behaving well.

  12485   Mon Sep 12 20:19:25 2016 LydiaUpdateGeneralMC REFL beam splitter not replaced

The beam splitter that directs light into the MC REFL photodiode has not been replaced; there is still a mirror there. Gautam suggested we wait to replace it until the PSL shutter is open so the beam can be aligned. However, this must be done before going to high power.

GV addendum: What I suggested was to try and recover the arm alignment using the current low power configuration after pumpdown - since we were well aligned just before pumpdown, we should be able to recover this alignment pretty easily at low power. After locking both arms and running the dither align (also center all Oplevs), we can go ahead do the following:

  • Replace mirror in MC Refl path with 10% reflection BS (Johannes, Lydia and I confirmed that this is on the AP table earlier today). Then align the reflected beam onto the PD using the tiny mirror
  • Replace HR mirror in Transmon path at the EY table
  • Replace ND filters on Transmon QPDs at EX and EY tables
  • Repalce ND filter on Transmon CCD at EY table
  • Revert MC autolocker to the nominal version instead of the low power version we have been using during the vent
  • Turn up MC to nominal power by rotating the wave plate on the PSL table - confirm that we have nominal levels by measuring with power meter
  • Recover single arm locks, green beatnotes etc at nominal operating conditions
     
  12484   Mon Sep 12 20:15:22 2016 LydiaUpdateSUSDiagonalization in air

[Lydia, Teng]

We ran the scripts to diagonalize the damping matrices using the free swinging data from staurday night/sunday morning. The actual entries used for damping have not been changed. However, we did generate updated matrices for all the main optics (not including the mode cleaner optics, which were not free swinging over the weekend).

  • The scripts appear to be mostly working as intended, with a couple of issues:
    • The plots made by makeSUSSpectra claim to be showing spectra of the individual OSEM readings, but are actually dofs calculated using the ideal input matrix.
    • The existing parameters file (for the peak finding) was only fitting the lorentz peaks to a very narrow band of data, close to the bandwidth of the spectrum. Too narrow a band means that the initial guess must be very close, and also means there are not enough points to fit to.
      • We modified a copy of the paramters file to use a wider band (~.1 Hz) for fitting, and also use updated estimates of the mode frequencies.
      • This was largely successful, but the ITMY POS peak is very close to the SIDE peak, and POS is also stringly coupled to SIDE, so the wider bandwidth fitting can't separate the peaks. (See attachment 1)
      • A longer time series, plus more accurate initial guesses for the resonance frequencies, would allow us to fit to a smaller (~.03 Hz) band without encountering the stated issues.
      • A better way than manually examining plots to choose an initial frequency guess would be to automatically start at the overall maximum point in the spectrum between 0.4 and 1.5 Hz
  • Most of the diagonalization results seem good: "Badness" numbers of 4-6 and secondary peaks very supressed or absent on spectra plotted in dof basis (See attachment 2). ITMY, perhaps beacuse of a related issue, has phase problems with the matrix elements that result in messages like "osem/dof 2/1 is imaginary."
Attachment 1: ITMY_fit.jpg
ITMY_fit.jpg
Attachment 2: BS_diag.jpg
BS_diag.jpg
  12483   Mon Sep 12 17:54:24 2016 SteveUpdateVACpumpdown 79

 We are pumping down. The annuloses are below 10 mTorr

The vacuum gauges are not communicating with the medm so there is no plot available.

The main volume pumping is stopped at P1 = 220 Torr  ( New SuperBee 174 Torr ) for overnight.

note: SuperBee was reading 791 Torr at atm and it was not set to 760 !

  12482   Mon Sep 12 17:15:22 2016 JohannesUpdateGeneralPRM SRM alignment

[Gautam, Steve, Johannes]

We put on the remaining heavy doors on the chambers (ITMY, ITMX,ETMX, in this order) this morning. On the ITMY and ETMX tables we placed old OpLev steering mirrors that are clean and baked as witness plates such that may one day provide some insight into dust accumulation on optics.

With the heavy doors on we confirmed that we were still able to lock both IFO arms and used the dither scripts to optimize the alignment. Following that we centered all OpLevs and aligned the X and Y green beams.

ITMY table witness plate
ETMX table witness plate

 

  12481   Sat Sep 10 14:26:01 2016 ranaUpdateGeneralPRM SRM alignment

For the PRM, I aligned it until the arm flashes were maximized and the REFL camera showed a centered spot with dips happening during the arm pops. AS port was more messy since the Michelson alignment wasn't perfect, but the spots were both near the center of the cam and the SRM alignment maximized the wangy fringiness of the image as well as the angry cat meow sounds that the full IFO makes as heard through the DAFI (listening to POX).

On Monday, Osamu should be back and can help with doors and then alignment recovery and locking.

Attachment 1: Osamu.JPG
Osamu.JPG
  12480   Fri Sep 9 17:50:02 2016 gautamUpdateSUSHeavy doors on BS-PRM, ETMY chambers

[steve, teng, johannes, lydia, gautam]

  • we set about doing some final checks on the Y arm while Johannes and Lydia worked on the X arm alignment
  • locked IMC, turned on Oplev HeNes for ITMY, SRM, PRM, BS and ETMY
  • I first went into the BS-PRM chamber. Traced Oplev paths for PRM and BS, checked that the beam is approximately centered on all the steering mirrors, and traced the beam with a clean beam card to make sure there was no clipping. The beams make it out of the vacuum onto the PDs, but are not centered
  • I also checked the Y arm green - the beam isn't quite centered on the periscope mirrors but I guess this has always been the case and I didn't venture to make any changes
  • Checked new PRM foil hats were secure
  • Checked the main IR beam out of the IMC, and also the IPANG beam - Steve suggested we keep track of the way this moves during pumpdown. However, I didn't quite think this through and we put the heavy door on the BS-PRM chamber before checking where the IPANG beam was on ETMY table (we later found that the beam was a tad too high. Anyways, this isn't critical, wouldve been nice to have this reference though
  • Checked that there were no tools lying around inside the chamber, and proceeded to put the heavy door on
  • Moved to ETMY table, and did much of the same as above - Oplev beam makes it successfully out off the ETM, OSEM cables aren't a risk to clipping the green input beam
  • Proceeded to put the heavy door on ETMY chamber
  • I would have liked to put the heavy door on the ITMY chamber today evening too, but while freeing the SRM from its EQ stops, I noticed that the LL and LR OSEM PD readouts are approximately 60 and 75 % of their saturation values. I think this warrants fixing (I also checked against the frame files from our last DRFPMI lock in march and the PD signals are significantly different) so we should do this before putting the heavy door on. It would also be a good idea to check the table leveling
  • The Oplev beams for ITMY and SRM make it cleanly out of the chamber so all looks good on that front
  • IR and green beams are well clear of any OSEM cables

Depending on how the X arm situation is, we will finish putting back all the heavy doors on Monday and start the pumpdown

GV Edit 11.30pm: 

  • We succeeded in locking the X arm as well, although the transmission peaked at 0.1 (but this is the high gain PD and not the QPD, and also, unlike the Y arm, the 50-50 BS splitting the transmitted light between the QPD and the high gain PD is still in place, so can't really compare with the Y arm value of 0.6)
  • To get the lock going, we had to change a bunch of things like the POX DC offset, demod phase, sign of the gain etc. It is unclear whether we are locking on the TEM00 mode, but we judged it is sufficient to close doors and pump down
  • Johannes and I centered the ETMX and ITMX OL spots on their respective QPDs. Earlier today, Johannes and Lydia had checked ITMX and ETMX OL paths, everything looks decent
  • JE piggyback edit : We also tied the upper ITMX OSEM cables to the suspension cage side using copper wire since particularly UR looked like it could slip and possibly fall down into the beam path
  • JE piggyback edit: While leveling the ITMX table, Gautam and I found that some of the screws that secure the weights were not vented. None of these were put in during this vent. We replaced them all with vented screws.
  • Rana also checked PRM and SRM alignment, all looks okay on that front - the OSEM problem I had alluded to earlier isn't really a problem, once the SRM is aligned, all the OSEMs are reasonably close to 50% of their saturation value.

Looks like on Monday, we will look to put the heavy doors on ITMY, ITMX and ETMX chambers, and begin the pumpdown

  12479   Fri Sep 9 11:24:15 2016 SteveUpdatesafetysafety training

Visiting graduate student Teng Zhang from Glasglow received 40m specific safety training yesterday.

  12478   Thu Sep 8 22:12:36 2016 gautamUpdateSUSX arm in place, locked to green, IR flashes visible

Detailed elog to follow but summary of todays activities:

  • ITMX and ETMX are back in their respective positions
  • F.C was peeled, OSEMs were inserted after releasing EQ stops
  • X arm was aligned to green
  • IMC was locked, BS was used to adjust IR input pointing till beam was cleanly passing through irides (slid on to the tower)
  • After best efforts for today - we see flashes as judged from TRX signal and also POX11_I. Unfortunately these are really weak and we can't lock, let alone see anything on the screens. Tomorrow we can try some more fine alignment
  12477   Wed Sep 7 18:00:47 2016 LydiaUpdateGeneralAcromag Progress

[Teng, Lydia]

We would like to establish a system for setting up ADC channels and integrating them into the existing EPICS framework, so that we can gradually switch over channels that are currently handled by the aging slow machines. Otherwise, we will be stuck when they eventually fail. As a preliminary test for this method, we are in the process of setting up an Acromag ADC to read the "Diagnostic" output of the PSL controller. This information will also be useful to monitor the health of the PSL. 

Today, we accomplished the following:

  • Configured the Acromag XT1221 for use on the martian netowrk. It is assigned the static IP 192.168.113.237, with hostname iocPSLmon. 
  • Connected the Acromag to a switch on the 1X6 cabinet, and set it up on a desk near the X arm door with a 24V DC power supply. 
  • Verified that the IP was reachable from a control room desktop. 
  • Modified the files from Aiden's wiki page (myiocconfig.cmd and IOCTEST.db) to reflect our setup. 
  • Attached input 0 to a DC voltage and retrieved the output over the network. 
    • Channel name: "C3:ACROMAG_INPUT0" 
    • Values are currently uncalibrated, the voltage is represented by a 16 bit signed integer
    • We changed the value of the DC input and verified that the channel output changed in the expected direction

The power supply has been turned off for the night. 

  12476   Wed Sep 7 17:23:26 2016 gautamUpdateSUSPRM LR fixed for now

[johannes, gautam]

  • We took the heavy door off in the morning with Steve's help
  • The problem was quickly identified as the Al foil on the back of the PRM OSEMs (placed to mitigate scattered light making it into the OSEM that was making locking difficult) shorting out the pins on the rear of the OSEM
  • We decided against using a black glass beam stop behind the PRM - rather, we decided to go for Al foil hats that were
    1. More "domed" - so the back plane of the OSEM isn't in direct contact with the Al foil, though the hats themselves are secure and shouldn't simply fly off during pump down etc
    2. Have a piece of kapton (courtesy Koji from the OMC lab) in the dome so that even if the foil hats move around slightly, there should be no danger of accidentally shorting out any pins
  • Without removing the PRM OSEMs, we were only able to image UR and UL unambiguously showing that they have no filters. Not a single of the 5 spare 'short' OSEMs have filters. We have to open the ITMY chamber to reposition the OSEMs in the near future, which is when we will inspect SRM for filters.
  • Attachment #1 shows a picture of these foil hats - the ones actually put on are shaped slightly differently, but the idea is the same
  • Attachment #2 shows the PRM with its new OSEM hats (we also used a piece of clean copper wire to tie the OSEM cables to the tower on the bottom left of the cage as viewed from the BS-PRM chamber door)
  • After closing up the BS-PRM chamber, I locked the IMC to see if the input pointing had gone way off because of our work on the table and the reputation of the tip-tilts hysteresis - I can see weak flashes in the Y arm but not enough to lock - so I will tweak the alignment a little
  • Once I can recover the Y arm alignment, we can move on to peeling first contact and putting the X arm optics in.

Edit 7.30pm: I have managed to recover Y-arm in air locking, and the transmission is up at ~0.6 again which is what we were seeing prior to touching anything on the BS-PRM table, so it looks like the tip-tilt has not gone badly astray... I have also restored the Satellite boxes so that both PRM and SRM have their designated boxes

Attachment 1: IMG_3208.JPG
IMG_3208.JPG
Attachment 2: IMG_3211.JPG
IMG_3211.JPG
  12475   Tue Sep 6 20:52:42 2016 gautamUpdateSUSOSEM issues

The modes look like they're at the right frequencies, so pointing more and more towards a LED or satellite box issue.

We peeked into the BS-PRM chamber via the ITMX chamber to see if we could shed any light on this situation. It's hard to get a picture that is in focus, but it looks quite clear that the LR LED (in the lower left when viewed from the HR side) isn't anywhere near as bright as the rest (see Attachment #1). Various hypothesis include failed LED / piece of Al foil blocking the LED / teflon aperture slipped over the LED. But looks like we can't solve this without opening up the BS-PRM chamber. The plan tomorrow is to open up the chamber, pull out the problematic coil. Once we have a better idea of what is going wrong, we can decide what the appropriate course of action is - replace the OSEM or something else. 


As part of the diagnosis, I switched the PRM and SRM satellite boxes earlier today evening around 6pm. They remain in this switched state for now.


Steve, we plan to take the BS-PRM heavy door off tomorrow morning.

Attachment 1: P9060254.JPG
P9060254.JPG
  12474   Tue Sep 6 20:45:14 2016 gautamUpdateSUSX arm test masses back in chamber

[Teng, Johannes, Lydia, gautam]

  • The goal was to peel F.C. off both the X arm test masses and start work on aligning the arm
  • However the F.C. peeling wasn't successful - Johannes spotted spme residual junk close to the center of the optic on ITMX and I saw a whole bunch of specks in and around the center of the ETM (see Attachment #1)
  • Moreover, the PRM LR OSEM issue meant that we decided to re-paint the X arm optics and only take it off after debugging this OSEM PD issue
  • Attachment #2 and #3 show the AR and HR face of the ITM respectively after F.C painting
  • Attachment #4 shows the ETM HR face after HR painting
  • Both towers have been moved, so any pre-emptive levelling has probably gone out the window, just something to be aware of when we put the towers back in place....
  • There looks to be some filaments of F.C towards the edge of both the ITM and the ETM. These have been successfully removed with isopropanol + optical tissue, we should take care to do so before peeling the F.C....
Attachment 1: IMG_3137.JPG
IMG_3137.JPG
Attachment 2: IMG_3143.JPG
IMG_3143.JPG
Attachment 3: IMG_3142.JPG
IMG_3142.JPG
Attachment 4: IMG_3148.JPG
IMG_3148.JPG
  12473   Tue Sep 6 20:30:56 2016 ranaUpdateSUSITMY_UL is sick

In the morning, Steve will start opening the north BS door so that we can enter to inspect the PRM LR OSEM.

For the ITMY, I squished together the cables which are in the 'Cable Interface Board' which lives in the rack. This thing takes the 64 pin IDC from the satellite module and converts it into 2 D-sub connectors to go to the PD whitening board and the coil driver board. Lets see if the ITMY OSEM glitches change character overnight.

  12472   Tue Sep 6 18:21:13 2016 ranaUpdateSUSOSEM issues

I looked at the PRM free swing spectra. The modes look like they're at the right frequencies, so pointing more and more towards a LED or satellite box issue.

  7/2011 9/2016
POS 0.993 0.94
PIT 0.612 0.60
YAW 0.833 0.76
SIDE 0.999 0.993

Some of the frequencies have changed between the 2011 in-vac measurement and our 2016 in-air measurement, but that seems within usual parameters.

Attachment 1: PRMswing.png
PRMswing.png
  12471   Tue Sep 6 00:14:14 2016 gautamUpdateElectronicsSatellite Amplifier

 

If we have some data with one of the optics clamped and the open light hitting the PD, or with the OSEMs removed and sitting on the table, that would be useful for evaluating the end-to-end noise of the OSEM circuit. It seems like we probably have that due to the vent work, so please post the times here if you have them.

The ETMX OSEMs have been attached to its Satellite box and plugged in for the last 10 days or so, with the PD exposed to the unobstructed LED. I pulled the spectrum of one of the sensors (mean detrended, I assume this takes care of removing the DC value?). The DQed channels claim to record um (the raw ADC counts are multiplied by a conversion factor of 0.36). For comparison, re-converted the y-axis for the measured curve to counts, and multiplied the total noise curve from the LISO simulation by a factor of 3267.8cts/V (2^16cts/20V) so the Y axis is noise in units of counts/rtHz. At 1Hz, there is more than an order of magnitude difference between the simulation and the measurement which makes me suspect my y-axis conversion, but I think I've done this correctly. Can such a large discrepancy be solely due to thick film resistors?

Attachment 1: osempdComparison.pdf
osempdComparison.pdf
  12470   Mon Sep 5 21:05:06 2016 ranaUpdateElectronicsSatellite Amplifier

The attached PDF shows the output noise of the satellite amp. This was calculated using 'osempd.fil' in the 40m/LISO GitLab repo.

The mean voltage output is ~1 Vdc, which corresponds to a current with a shot noise level of 100 nV/rHz on this plot. So the opamp current noise dominates below 1 Hz as long as the OSEM LED output is indeed quantum limited down to 0.1 Hz. Sounds highly implausible.

To convert into meters, we divide by the OSEM conversion factor of ~1.6 V/mm, so the shot noise equivalent would be ~1e-10 m/rHz above 1 Hz.

After adding the sat amp to the 40m DCC tree (D1600348), I notice that not only is the PD readout not built for low noise, neither is the LED drive.  The noise should be dominated by the voltage noise of the LT1031 voltage reference. This has a noise of ~500 nV/rHz at 1 Hz. That corresponds to an equivalent current noise through the LED of 25 mA * (500e-9 / 10) ~ 1 nA/rHz. Or ~45 nV/rHz at the sat amp output. This would be OK as long as everything behaves ideally. BUT, we have thick film (i.e. black surface mount) resistors on the LED drive so we'll have to measure it to make sure.

Also, why is the OSEM LED included in the feedback loop of the driver? It means disconnecting the cable from the sat amp makes the driver go unstable probably. I think one concept is that including the device in the feedback loop makes it so that any EMI picked up in the cabling, etc. gets cancelled out by the opamp. But this then requires that we test each driver to make sure it doesn't oscillate when driving the long cable. frown 

If we have some data with one of the optics clamped and the open light hitting the PD, or with the OSEMs removed and sitting on the table, that would be useful for evaluating the end-to-end noise of the OSEM circuit. It seems like we probably have that due to the vent work, so please post the times here if you have them.

Attachment 1: osempd.pdf
osempd.pdf
  12469   Mon Sep 5 19:57:24 2016 OsamuUpdateSUSOSEM adjustments

Hi 40m people,

As Rana is saying, the bounce mode does not matter, or we cannot do anything. Generally speaking, the bounce mode cannot be damped by the setting of 40m SUS. Some tweak techniques may damp a bounce mode by res-gain or something, but it is not a proper way, I think.

This is also that Rana is already saying that the important thing is to find a good direction of OSEM to hit the LED beam to the magnet. Even if the magnet is not located at the center of OSEM hole, still you can find the optimal orientation of OSEM to hit the LED beam to the center of magnet by rotating the OSEM.

I know only an old document of T040054 that Shihori summarized how to adjust the matrix at the 40m. Too bad input/output matrix may introduce some troubles, but even roughly adjusted matrix should be still fine.

I will be at Caltech on 12-14 of September. If I can help something, I am willing to work with you!

  12468   Fri Sep 2 21:16:45 2016 ranaUpdateElectronicsSatellite Amplifier

In November of 2010, Valera Frolov (LLO), investigated our satellite amplifiers and made some recommendations about how to increase the SNR.

In light of the recent issues, we ought to fix up one of the spares into this state and swap it in for the ITMY's funky box.

The sat amp schematic is (D961289). It has several versions. Our spare is labeled as version D (not a choice on the DCC page).

  1. U13-17 seem superfluous. Why would we need a high current buffer to drive the slow EPICS ADCs ?
  2. The Radd resistors indicated on the B2 schematic have not been added to this board. What would be the purpose of adding them anyway?
  3. The transimpedance resistors are, in fact, 29.4k as indicated on the B1 schematic.
  4. The LT1125: V_n = 3 nV w/ f_corner = 5 Hz. I_n = 0.4 pA w/ f_corner = 100 Hz. So at 1 Hz the current noise is limiting at ~120 nV/rHz or 650 nV/rHz using the 161k that Valera recommends.
  5. Even in that case the ADC noise is higher than the opamp noise.
  6. We should be using metal film resistors instead of the thick film junk that's installed now.

Edit (Sep 6): The purpose of the Radd resistors is to lower the resistance and thus up the current through the LED. The equivalent load becomes 287 Ohms. Presumably, this in series with the LED is what gives the 25 mA stated on the schematic. This implies the LED has an effective resistance of 100 Ohms at this operating point. Why 3 resistors? To distribute the heat load. The 1206 SMD resistors are usually rated for 1/4 W. Better to replace with 287 Ohm metal film resistors rated for 1 W, if Steve can find them online.

  12467   Fri Sep 2 21:15:53 2016 gautamUpdateSUSITMX and ETMX preemptive table leveling

The ITMX table had relaxed overnight into a slightly misaligned state overnight - since the ITMX table holds PR2 and hence can affect the input pointing, we decided to fix this before commencing alignment work today. The misalignment was not as bad as what Johannes observed prior to his first re-leveling attempt, but was ~1 division on the spirit level. So I decided to move one set of weights to level the table again. It is entirely possible that over the next couple of days, the table will shift slightly again, but the hope is that we are closer to the 'ideal' orientation of the table now... Pictures to follow...

ELOG V3.1.3-